OpenStudy (anonymous):

A parachutist with a camera, both descending at a speed of 20 m/s, releases that camera at an altitude of 140 m. What is the velocity of the camera just before it hits the ground?

OpenStudy (gebooors):

Acceleration of the camera is g after releasing. It hits ground before Parachutist

OpenStudy (anonymous):

No, wrong, you need velocity, not acceleration, so use y=yo + vt + 1/2at^2

OpenStudy (anonymous):

I used the equation vf^2 = vi^2 + 2at I got 48.41 but apparently that's wrong

OpenStudy (anonymous):

cause you need an equation implementing time.... though I am not sure about the time..check on the internet :)

OpenStudy (anonymous):

what does the equation implementing time mean? I know the units are m/s since im finding final velocity

OpenStudy (anonymous):

yes, but you need to use the equation I did... with time... then you can find from acceleration to time, you know the aceleration has to be 9.8m/s^2... but then the velocity is found with that altidude they have given...

OpenStudy (anonymous):

y=yo + vt + 1/2at^2 is the equation you gave me. what is the y and the yo?

OpenStudy (anonymous):

y = final altidude and yo is initial altiditue

OpenStudy (anonymous):

does that just mean final and initial velocity?

OpenStudy (anonymous):

no, displacment, like the height

OpenStudy (anonymous):

so y= 140 + 20 + 1/2 * -9.8 * t^2